Suppose the demand for a product is given by Qc=200 - 2Pt+Pc where Qc is the quantity demand for coffee,Pc is the price of coffee,Pt is the price of the substitute good tea.The price of the substitute good is  ₹400/kg.Suppos Pc=₹650/kg,what is the price elasticity of demand?


1
Expert's answer
2021-02-23T13:26:18-0500
"p_t=400"

"p_c=650"

"Q=200-2\times400+650=50"

"E=1\times \frac {650}{50}=13"